# CasTGAN: Cascaded Generative Adversarial Network for Realistic Tabular Data Synthesis

### v1

@@ 4.4 Generators and Discriminator
5 Experimental Setup
Evaluating the performance of GANs is a non-trivial task, and this can be evident from the variety of GAN evaluation criteria in literature. In this section, we describe the experimental design for CasTGAN, with a thorough discussion of the metrics used for the model’s analysis. We implemented our model using PyTorch in Python on a Linux Ubuntu 20.04 machine running on AMD Ryzen Threadripper 3990X and Nvidia GeForce RTX 3090. Our CasTGAN source code is publicly available [^1] https://github.com/abedshantti/CasTGAN.
5.1 Datasets
Our CasTGAN is designed for synthesis of tabular data that can be typically found in the financial and healthcare sectors. We therefore use tabular mixed datasets characterised by a combination of categorical and numerical features. We use four datasets where the task is the binary classification of the target label - Adult [21], Bank Marketing [27], Taiwan Credit [44] and Diabetes [36]. Meanwhile, we use the House Prices [19] and Cars [37] datasets for regression. We additionally highlight that binary columns in the datasets are handled as categorical variables. An overview of the datasets used is presented in Table 1. For synthesising data with CasTGAN and other baselines, we use 50% of the datasets’ total number of samples for training the models. The remaining 50% of the data samples is dedicated for evaluating the generated synthetic output.
Table 1: Datasets used in this study.
@@ 5.2 Baselines
5.3 Evaluation Criteria
5.3.1 Train on Synthetic, Test on Real (TSTR)
@@
We measure reliability of the synthetic output produced by CasTGAN by training machine learning models on the generated data. We fitted three machine learning models on the generated output - namely AdaBoost, random forest and logistic regression for classification tasks and AdaBoost, decision trees and Linear SVM for regression tasks. We then used the trained models to predict the target label of the test data and we report the precision-recall area under curve (PR-AUC) for binary classification tasks and the R2 score for regression tasks. Since our main objective is to measure the machine learning utility of the generated data rather than assessing the individual performance of each classifier, we average the metrics produced by the three machine learning models.
5.3.2 Univariate Distributions
@@
We also assess how well do the individual features generated by CasTGAN resemble the features of the real data. As we quantitatively analyse how well our model learns the univariate feature distributions, we first one-hot encode and normalise the synthetic and real datasets. We calculate the dimension-wise mean of the individual features of the synthetic output and training data and report the RMSE between the real and the synthetic output. Additionally, we report the Kolmogorov-Smirnov two-sample test score [4] between the real univariate features and the synthetic ones.
5.3.3 Correlation and Diversity
@@
Measuring the validity of the synthetic output represents a significant challenge for GAN frameworks. In computer vision applications of GANs, the synthetic images can be in some cases distinguished from the real images by a human observation of irregularities in the output such as pupil orientation in human eyes and out-of-place pixels. In tabular GANs, similar challenges exist as there is no standard approach in the literature for quantifying the proportion of invalid samples in the synthetic output. For instance, a common example in the tabular synthesis literature is highlighting how an entry such as gender = "Female" and diagnosis = "Prostate Cancer" is by definition an invalid record, as no such entry exists in the original data, nor can a female be diagnosed with prostate cancer by a physician. In financial datasets, such nested relationships also exist between features if for example we inspect a dataset with a city column and a country column. In such cases, a record with city = "Buenos Aires" and country = "Malta" is by definition invalid. Meanwhile, entries with city = "Alexandria" and country = "United States" is valid as Alexandria exists in the United States even though it is more commonly attributed to the country "Egypt". It is for this reason that quantifying the invalid output generated by tabular GANs is no easy task, even in the presence of domain knowledge.
@@
Given that public datasets are used in this work for the purpose of reproducability, we do not have the full domain knowledge for these datasets, thus, we propose an alternative method for quantifying the validity of the synthetic data. First, we consider calculating the difference in feature correlations between the training data and the fake data. For computing the correlations between numerical features we use the Pearson’s correlation coefficient, while the Cramer’s V measure is used for capturing the correlation between categorical features. The correlation score is found by calculating the RMSE score between the elements of the triangular matrix of the synthetic dataset and the real dataset.
For measuring the diversity of the categorical output, we propose a new metric - Unique Pairwise Categorical Combinations (UPCC). In essence, we count the total number of unique interactions between any pair of categorical features in the dataset. For instance, in the Adults dataset the combinations [education = "Bachelors" and marital-status = "Never-married"], [education = "Bachelors" and sex = "Male"] and [marital-status = "Never-married" and sex = "Male"] each counts as a single pairwise combination, regardless of how many times they appear in the data. A reliable model therefore ensures that the UPCC of its output should be comparable to that of the original data. Subsequently, the UPCC Ratio is the number of unique pairwise combinations of synthetic output divided by the total number of unique combinations of the training data.
Finally, we estimate the validity of the model’s output by dividing the correlation RMSE score of the model by the UPCC Ratio. We name this measure as the CORDV score. A lower CORDV score indicates that the model is able to minimise the difference in feature correlation between its synthetic output and the real data, while simultaneously not impeding its ability in generating unique feature combinations. Meanwhile, a worse generative model is reflected a greater CORDV score indicating that the model poorly captures the correlations while potentially restricting the uniqueness of the categorical pairs.
5.3.4 White-box privacy attacks
@@
6 Results
6.1 Machine Learning Utility
We use the synthetic data generated by CasTGAN for fitting machine learning classification and regression supervised models on the six datasets and compare our performance on the test set against models fitted on the training set and models fitted on the five synthetic output of the five baseline methods. The results are computed in Table 2.
Table 2: Binary classification (PR-AUC score) and regression (R2 score) evaluation on the test sets.
@@ 6.1 Machine Learning Utility
From Table 2 we can observe that the TSTR metrics for our CasTGAN is consistently within the best performing synthetic output, outperforming all the baselines on three out of six datasets. Meanwhile, we note that medGAN demonstrated the best results on regression datasets. In addition, CTGAN and CTAB-GAN have comparatively close performance to our method and the training examples. We note that we were unable to run the highly-dimensional Cars dataset on CTAB-GAN to exceedingly large memory requirements attributed to representing wide datasets in convolutional neural network GAN-based approaches.
6.2 Univariate Similarity
To quantitatively analyse how well CasTGAN represents the feature distributions of the original data, we compare the Euclidean distance RMSE and Kolmogorov-Smirnov statistic between the synthetic data and the real data in Table 3. Generally, we note that our CasTGAN, medGAN and CTAB-GAN dominate the dimension-wise statistical similarity test. It can be observed that CasTGAN represents the features of the Adult and Cars datasets particularly well, while performing comparatively on the remaining datasets. We can therefore deduce that CasTGAN is particularly useful for datasets with a greater number of unique categories.
Table 3: Univariate dimension-wise statistical comparison.
@@ 6.2 Univariate Similarity
6.3 Output Validity
We quantify the validity of the output by considering the number of unique pairwise categorical combinations (UPCC), the correlation error between the synthetic and the real data, and the CORDV score, which is essentially the correlation divided by the UPCC ratio. From Table 4, we can observe how the different generative methods rank among the three aforementioned metrics. First, we note that CTGAN and medGAN perform well in generating a large number of unique feature combination, in most cases, even more than the number of combinations that can be found in the training data. This is particularly impressive for the Cars dataset, where both models managed to generate more than twice the number of categorical combinations of the training set. Similarly, we also observe that CTAB-GAN performs relatively well in exploring diverse categorical combinations. Meanwhile, we notice that our CasTGAN is more conservative when it comes to generate unique categorical combinations. For all the dataset, CasTGAN produces a marginally lower number of pairwise combinations than can be typically found in the training set. Moreover, the UPCC can be a good indicator of mode collapse and this is reflected by the significantly low UPCC values for table-gan and cWGAN, where it can be deduced that these models generated a limited number of modes for some categories.
Table 4: Diversity and correlation comparison.
@@ 6.3 Output Validity
In contrast, it can be observed from Table 4 that CasTGAN generally outperforms the other baselines in capturing the feature correlations of the datasets. The lower correlation RMSE score entails that CasTGAN prioritises the representation of correlations and feature interdependence in the real data. Meanwhile, the CORDV score aims to quantify the trade-off between the diversity and the proximity of the synthetic data to the real data. We observe from the results that the highest CORDV scores are evenly split among our CasTGAN and CTGAN.
@@
  • Adult dataset: we use the fields "relationship" and "sex" to calculate the number of invalid records. We posit that if "relationship" = "Husband", then the sex feature needs to be set to "Male". Likewise, the "relationship" = "Wife" needs to align the gender field assigned as "Female". This is not based on our assumptions, but rather running explanatory data analysis on the training set confirms that the records are matched in such a manner.
  • Cars dataset: we use the fields "make" and "model" and classify a synthetic sample as invalid if the synthetic car’s "model" does not in fact belong to the "make" that can be found in training data.
  • Housing dataset: the housing dataset consists of the fields "year built" and "year renovation". Logically, a property cannot be renovated before it was built, and further inspecting the data indeed confirms that there are no observations with "year renovation" before "year built".
Based on the aforementioned fronts, the ratio of invalid records generated by CasTGAN and three baseline methods are demonstrated in Table 5. We chose not to include table-GAN and cWGAN in the comparison as the synthetic output is characterised by major mode collapse.
Table 5: The ratio of invalid synthetic records.
@@
medGAN 0.2905 0.9111 0.5929
CTGAN 0.0982 0.4289 0.5475
CTAB-GAN 0.0414 - 0.4442
CasTGAN 0.0067 0.2782 0.2298
From Table 5 we can observe that CasTGAN remarkably reduced the number of invalid synthetic records of the Adult dataset. Our method also significantly decreased the number of invalid records in the Cars dataset. This resembles a major improvement from CTGAN, while noting the challenging nature of modelling the Cars dataset due to the large number of categories present. Furthermore, it is evident that CasTGAN outperforms the other generative approaches on the numerical features of the Housing dataset.
6.4 Robustness Against Privacy Attacks
For conducting white-box privacy attacks, we set the number of attacking iterations to five, where each feature in the synthetic data is updated five times based on the output of the auxiliary learners. Moreover, the membership attacks are launched on 10% of the total number of overall samples. We highlight that the ratio of attacked samples does not impact the evaluation of the robustness of our approach as we only compute the attack distance metrics with respect to the attacked samples. The Euclidean distance of the attacked samples to the training data and to the synthetic data prior to the membership attacking is computed in Table 6.
Table 6: White box privacy attacks on the auxiliary learners proximity measures.

### v2

@@ IV-D Generators and Discriminator
V Experimental Setup
Evaluating the performance of GANs is a non-trivial task, and this is evident from literature, where no standard approach for evaluating GANs on tabular data can be found. In this section, we describe the experimental design for CasTGAN, with a thorough discussion of the metrics used for the model’s analysis. We implemented our model using PyTorch in Python on a Linux Ubuntu 20.04 machine running on AMD Ryzen Threadripper 3990X and Nvidia GeForce RTX 3090. Our CasTGAN source code is publicly available [^1] https://github.com/abedshantti/CasTGAN.
V-A Datasets
Our CasTGAN is designed for synthesis of tabular data that can be typically found in the financial and healthcare sectors. We therefore use tabular mixed datasets characterised by a combination of categorical and numerical features. We use four datasets where the task is the binary classification of the target label - Adult [47], Bank Marketing [48], Taiwan Credit [49] and Diabetes [50]. Meanwhile, we use the House Prices [51] and Cars [52] datasets for regression. We additionally highlight that binary columns in the datasets are handled as categorical variables. An overview of the datasets used is presented in Table II. For synthesising data with CasTGAN and other baselines, we use 50% of the datasets’ total number of samples for training the models. The remaining 50% of the data samples is dedicated for evaluating the generated synthetic output.
TABLE II: Datasets used in this study.
@@ V-B Baselines
V-C Hyperparameter selection
We emphasise that while our framework comprises an ample number of hyperparameters, we state that our reasoning for refraining from fine-tuning our model’s hyperparameters in our experimental results is twofold. First, an extensive hyperparameter tuning would need to be conducted on a data-level basis and on a criteria level. To demonstrate that our framework is compatible with any dataset, we conduct our experimental analysis and train CasTGAN on all the datasets in this work using the same set of parameters in Appendix A-A. The choice of these parameters is based on preliminary, yet limited experimentation of our framework which yielded satisfactory and promising synthetic output during the conceptualisation phase. From a criteria-level perspective, we highlight that the model settings that attain the best performance on an evaluation criterion do not necessarily improve the synthetic output on all performance evaluation fronts. For instance, there might exist a hyperparameter trade-off between maximising the machine learning utility and improving the univariate feature representation. Second, given that we refrain from conducting hyperparameter tuning for the benchmark models and instead adopt the recommended hyperparameters of the benchmarks across all datasets, the comparison between our method and the baselines can be regarded as fair and equitable, provided that no extensive tuning of CasTGAN is undertaken correspondingly. Notwithstanding, we refer the reader interested in strategies for selecting hyperparameters for training CasTGAN on their own data to Appendix A-B.
V-D Evaluation Criteria
V-D1 Train on Synthetic, Test on Real (TSTR)
@@
We measure reliability of the synthetic output produced by CasTGAN by training machine learning models on the generated data. We fitted three machine learning models on the generated output - namely AdaBoost, random forest and logistic regression for classification tasks and AdaBoost, decision trees and Linear SVM for regression tasks. We then used the trained models to predict the target label of the test data and we report the precision-recall area under curve (PR-AUC) for binary classification tasks and the $R^{2}$ score for regression tasks. Since our main objective is to measure the machine learning utility of the generated data rather than assessing the individual performance of each classifier, we average the metrics produced by the three machine learning models.
V-D2 Univariate Distributions
@@
We also assess the extent at which the individual features generated by CasTGAN resemble the features of the real data. As we quantitatively analyse how well our model learns the univariate feature distributions, we first one-hot encode and normalise the synthetic and real datasets. We calculate the dimension-wise mean of the individual features of the synthetic output and training data and report the RMSE between the real and the synthetic output. Additionally, we report the Kolmogorov-Smirnov two-sample test score [53] between the real univariate features and the synthetic ones.
V-D3 Correlation and Diversity
@@
Measuring the validity of the synthetic output represents a significant challenge for GAN frameworks. In computer vision applications of GANs, the synthetic images can be in some cases distinguished from the real images by a human observation of irregularities in the output such as pupil orientation in human eyes and out-of-place pixels. In tabular GANs, similar challenges exist as there is no standard approach in the literature for quantifying the proportion of invalid samples in the synthetic output. For instance, a common example in the tabular synthesis literature is highlighting how an entry such as gender = ”Female” and diagnosis = ”Prostate Cancer” is by definition an invalid record, as no such entry exists in the original data, nor can a female be diagnosed with prostate cancer by a physician. In financial datasets, such nested relationships also exist between features if for example we inspect a dataset with a city column and a country column. In such cases, a record with city = ”Buenos Aires” and country = ”Malta” is by definition invalid. Meanwhile, entries with city = ”Alexandria” and country = ”United States” is valid as Alexandria exists in the United States even though it is more commonly attributed to the country ”Egypt”. It is for this reason that quantifying the invalid output generated by tabular GANs is no easy task, even in the presence of domain knowledge.
@@
Given that public datasets are used in this work for the purpose of reproducability, we do not have the full domain knowledge for these datasets, thus, we propose an alternative method for quantifying the validity of the synthetic data. First, we consider calculating the difference in feature correlations between the training data and the fake data. For computing the correlations between numerical features we use the Pearson’s correlation coefficient, while the Cramer’s V measure is used for capturing the correlation between categorical features. The correlation score is found by calculating the root mean squared error (RMSE) score between the elements of the triangular matrix of the synthetic dataset and the real dataset.
For measuring the diversity of the categorical output, we propose a new metric - Unique Pairwise Categorical Combinations (UPCC). In essence, we count the total number of unique interactions between any pair of categorical features in the dataset. For instance, in the Adults dataset the combinations [education = ”Bachelors” and marital-status = ”Never-married”], [education = ”Bachelors” and sex = ”Male”] and [marital-status = ”Never-married” and sex = ”Male”] each counts as a single pairwise combination, regardless of how many times they appear in the data. A reliable model therefore ensures that the UPCC of its output should be comparable to that of the original data. Subsequently, the UPCC Ratio is the number of unique pairwise combinations of synthetic output divided by the total number of unique combinations of the training data.
Finally, we estimate the validity of the model’s output by dividing the correlation RMSE score of the model by the UPCC Ratio. We name this measure as the CORDV score. A lower CORDV score indicates that the model is able to minimise the difference in feature correlation between its synthetic output and the real data, while simultaneously not impeding its ability in generating unique feature combinations. Meanwhile, a worse generative model is reflected by a greater CORDV score, indicating that the model poorly captures the correlations while potentially restricting the uniqueness of the categorical pairs.
V-D4 White-Box Privacy Attacks
@@
VI Results
VI-A Machine Learning Utility
We use the synthetic data generated by CasTGAN for fitting machine learning classification and regression supervised models on the six datasets and compare our performance on the test set against models fitted on the training set and models fitted on the five synthetic output of the five baseline methods. Additionally, we compare the performance of our model and the other baselines against the real training datasets used for fitting the predictive models, which we refer to as Identity. We note that we were unable to run the highly-dimensional Cars dataset on CTAB-GAN to exceedingly large memory requirements attributed to representing very high-dimensional data in the convolutional neural network GAN-based approach. The results are computed in Table III.
TABLE III: Binary classification (PR-AUC score) and regression ($R^{2}$ score) evaluation on the test sets.
@@ VI-A Machine Learning Utility
From Table III we can observe that the TSTR metrics for our CasTGAN is consistently within the best performing synthetic output, outperforming all the baselines on three out of six datasets. On the Bank dataset, we can observe that CasTGAN’s PR-AUC score of 0.5657 ranks closer to the classification models trained on the Bank real training data with a score of 0.6085, than the second best classification by CTAB-GAN with a score 0.4920. Similarly, CasTGAN falls 0.0114 short of the PR-AUC score exhibited by the real data on the Diabetes dataset, whilst outperforming the second best synthetic model, CTGAN, with a PR-AUC difference of 0.0155. We also observe that CasTGAN also ranks first among the GAN models on the Credit dataset, despite being closely challenged by CTAB-GAN. For the datasets which CasTGAN did not achieve the highest machine learning utility, the results in Table III show that CasTGAN narrowly underperformed against two of the baselines on the Adult dataset. Furthermore, we note that medGAN demonstrated the best results on regression datasets, whereas, synthetic data from CasTGAN followed as the second best in terms of the $R^{2}$ score on the regression datasets. In general, the prediction results on the test sets suggest that synthetic data produced by CasTGAN is well suited to fitting machine learning models, as the predictive performance is comparable to training on the real data, and in-line with the best performing state-of-the-art tabular GANs.
VI-B Univariate Similarity
(a) Univariate categorical feature-wise distribution comparison between the real data (in blue) and the CasTGAN synthetic data (in orange). The counts of the categories are log-scaled for intuitively representing infrequent categories.
(b) Univariate numerical feature-wise distribution comparison between the real data (in blue) and the CasTGAN synthetic data (in orange). The distributions are plotted by applying Gaussian kernel density estimation.
Figure 4: Discrete and continuous univariate features distribution plots for the Bank dataset.
TABLE IV: Univariate dimension-wise statistical comparison showing the Euclidean root mean squared error (Euc. RMSE) and the Kolmogorov-Smirnov two-sample test score (KS statistic)
@@ VI-B Univariate Similarity
It is imperative that synthetic data generation techniques need to emulate the distribution of the features of the real data. One method for qualitatively evaluating the statistical similarity between the real data and the synthetic data is to visually compare the distributions for categorical and numerical attributes. Subsequently, we choose to display the comparison between the synthetic and real features for the Bank dataset, as it consists of a diverse and heterogeneous set of features. The depiction is demonstrated in Figure 4.
From Figure 4, we can observe how well CasTGAN performs in approximating the distributions for the categorical and numerical attributes. For categorical variables in Figure 3(a), it is evident that CasTGAN preserves the frequency of unique categories under each discrete variable. It can also be observed that our framework can successfully represent and sample the less frequent categories in the dataset. For the numerical features in Figure 3(b), we can find that the synthetic data distribution closely approximates the numerical distribution the real data. While adopting the Variational Gaussian Mixture models in CasTGAN helps in improving the density estimation of numerical features, as indicated by the number of peaks and the extensive coverage of the numerical bounds, we notice that approximation is slightly conservative, thus, does not represent extreme values in the real data’s numerical range to avoid violating boundary constraints and generating invalid records. The combined results of the numerical and categorical features distributions of the synthetic data from CasTGAN, in addition to the reasonable sampling of less frequent features, as demonstrated in Figure 4, signifies that our framework clearly does not suffer from mode collapse.
To quantitatively analyse how well CasTGAN represents the feature distributions of the original data, we compare the Euclidean distance RMSE and Kolmogorov-Smirnov statistic between the synthetic data and the real data in Table IV. We find that CasTGAN performs considerably the best in terms of the Euclidean distance RMSE on three of the six benchmark datasets: Adult, Bank, and Cars, while ranking comparatively to the best performing GANs on the three remaining datasets. It can also be observed that while CasTGAN only achieves the best KS statistic on the Cars dataset, and often trailing only slightly behind CTAB-GAN on the other datasets, the KS statistic by our model regularly maintains a small value, which can be interpreted as the higher likelihood of CasTGAN’s synthetic data to come from the same distribution as the real data. Generally, we note that our CasTGAN, along with medGAN and CTAB-GAN dominate the dimension-wise statistical similarity test. It is also apparent that CasTGAN represents the features of the Adult and Cars datasets particularly well, while performing comparatively on the remaining datasets. We can therefore deduce that CasTGAN is particularly useful for datasets with a greater number of unique categories.
VI-C Output Validity
In addition to ensuring univariate similarity, it is equally fundamental to evaluate how well the synthetic models preserve the interactions between the different features of a dataset. Inspecting the correlations between the data features of the real data and comparing the correlations with the synthetic data representation can indicate whether the generative data models are capable of simulating the relationship between the data variables from the real data representation. Figure 5 depicts the correlation matrices for the Adult real training data and the synthetic Adult data generated by CasTGAN. We can see from Figure 5 that our model notably learns the correlations between the real data features during the training process. By comparing Figure 4(a) and Figure 4(b), it is evident that there is only a marginal difference between the correlations of the real data and the CasTGAN synthetic data, the highest of which we note is a correlation difference of 0.14 between the ”native-country” and ”education-num” features. Otherwise, the model successfully emulates both the magnitude of the correlations and the sign, as to whether the feature correlations are positive or negative.
(a) Real data correlation matrix
(b) CasTGAN synthetic data correlation matrix
Figure 5: Adult dataset correlation map plots for the real and the synthetic data generated by CasTGAN. Larger absolute values indicate a stronger correlation, either positively or negatively.
TABLE V: Diversity and correlation comparison demonstrating the number of unique pairwise categorical combinations (UPCC), the correlation root mean squared error (Corr. RMSE) and the CORDV score.
@@ VI-C Output Validity
As emphasised in the motivation for this work, there is a need to establish evaluation frameworks that can critically assess the realistic-ness of the synthetically generated data. We therefore quantify the validity of the output by considering the number of unique pairwise categorical combinations (UPCC), the correlation error between the synthetic and the real data, and the CORDV score, which is essentially the correlation divided by the UPCC ratio. From Table V, we can observe how the different generative methods rank among the three aforementioned metrics. First, we note that CTGAN and medGAN perform well in generating a large number of unique feature combination, in most cases, even more than the number of combinations that can be found in the training data. This is particularly impressive for the Cars dataset, where both models managed to generate more than twice the number of categorical combinations of the training set. Similarly, we also observe that CTAB-GAN performs relatively well in exploring diverse categorical combinations. Meanwhile, we notice that our CasTGAN is more conservative when it comes to generate unique categorical combinations. For all the datasets, CasTGAN produces a marginally lower number of pairwise combinations than can be typically found in the training set. Moreover, the UPCC can be a good indicator of mode collapse and this is reflected by the significantly low UPCC values for table-gan and cWGAN, where it can be deduced that these models generated a limited number of modes for some categories.
In contrast, it can be observed from Table V that CasTGAN generally outperforms the other baselines in capturing the feature correlations of the datasets. The lower correlation RMSE score entails that CasTGAN prioritises the representation of correlations and feature interdependence in the real data. Meanwhile, the CORDV score aims to quantify the trade-off between the diversity and the proximity of the synthetic data to the real data. We observe from the results that the highest CORDV scores are evenly split among our CasTGAN and CTGAN.
@@
  • Adult dataset: we use the fields ”relationship” and ”sex” to calculate the number of invalid records. We posit that if ”relationship” = ”Husband”, then the sex feature needs to be set to ”Male”. Likewise, the ”relationship” = ”Wife” needs to align the gender field assigned as ”Female”. This is not based on our assumptions, but rather running exploratory data analysis on the training set confirms that the records are matched in such a manner.
  • Cars dataset: we use the fields ”make” and ”model” and classify a synthetic sample as invalid if the synthetic car’s ”model” does not in fact belong to the ”make” that can be found in training data.
  • Housing dataset: the housing dataset consists of the fields ”year built” and ”year renovation”. Logically, a property cannot be renovated before it was built, and further inspecting the data indeed confirms that there are no observations with ”year renovation” that precedes ”year built”.
Based on the aforementioned fronts, the ratio of invalid records generated by CasTGAN and three baseline methods are demonstrated in Table VI. We chose not to include table-GAN and cWGAN in the comparison as their synthetic output was found to be characterised by major mode collapse.
TABLE VI: An outline of the ratio of invalid synthetic records of the baseline GANs and our model.
@@
medGAN 29.05% 91.11% 59.29%
CTGAN 9.82% 42.89% 54.75%
CTAB-GAN 4.14% - 44.42%
CasTGAN 0.67% 27.82% 22.98%
From Table VI we can observe that CasTGAN remarkably reduced the number of invalid synthetic records of the Adult dataset. Our method also significantly decreased the number of invalid records in the Cars dataset. This resembles a major improvement from CTGAN, while noting the challenging nature of modelling the Cars dataset due to the large number of categories present. Furthermore, it is evident that CasTGAN outperforms the other generative approaches on the numerical features of the Housing dataset.
VI-D Robustness Against Privacy Attacks
For conducting white-box privacy attacks, we set the number of attacking iterations to five, where each feature in the synthetic data is updated five times based on the output of the auxiliary learners. Moreover, the membership attacks are launched on 10% of the total number of overall samples. We highlight that the ratio of attacked samples does not impact the evaluation of the robustness of our approach as we only compute the attack distance metrics with respect to the attacked samples. The Euclidean distance of the attacked samples to the training data and to the synthetic data prior to the membership attacking is computed in Table VII.
TABLE VII: White box privacy attacks on the auxiliary learners proximity measures.